Status Report

NASA Space Station On-Orbit Status 8 Aug 2003

By SpaceRef Editor
August 8, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 8 Aug 2003

All ISS systems continue to function nominally, except as noted previously or below.

The crew started another round of maintenance on the Russian segment (RS) ventilation system in the FGB (funktsionalnyi-grusovoi blok), with CDR Yuri Malenchenko first cleaning the removable screens of the three gas-liquid heat exchangers (GZhT), while FE/SO Ed Lu exchanged the two PS1/PS2 dust collector filters, disposing of the old filters (last time done: 7/11).  Yuri later went on to clean ventilation louvers on FGB panels as well as the protective mesh screens of the TsV ventilation fans (last time done: 7/10).

Ed Lu performed the regular once-a-week maintenance reboot on the operational PCS (portable computer system) laptops in the station.

Today’s schedule also called for another weekly inventory audit by the Science Officer of the available CWCs (contingency water containers) and their contents, to keep track of onboard water supplies.  [Last time done: 8/1.]

Working another session on the MSG (Microgravity Science Glovebox) with the CSLM-2 (Coarsening in Liquid/Solid mixtures #2) experiment, Ed Lu today supervised the termination of the SPU-6 (sample processing unit #6) sample processing after its 48-hr. “soaking” period, when the ECU (electronic control unit) sent a signal to the SPU to trigger the quench.  [Ed Lu then initiated transfer of the run data that were recorded by the ECU onto the MLC (MSG laptop computer), and finally removed and stowed SPU-6 as well as the remainder of the CSLM hardware.  The SAMS (space acceleration measurement system) was left in the work volume for use in the upcoming PFMI (Pore Formation and Mobility Investigation) operations (assuming SAMS can be restored to nominal operations).]

Yuri conducted the regular weekly checkout of the IP-1 air flow sensors which are installed in the various RS (Russian segment) hatchways and the hatchway between FGB and Node.

In the Service Module (SM), the CDR started another regeneration cycle on absorbent bed #1 of the BMP harmful impurities unit, leaving channel 2 in Purify mode.  [The “bakeout” cycle in the filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

Because of some safety concern on the ground, Ed Lu was asked to relocate CTB (cargo transfer bag) stowage from behind a Node panel (NOD1O1) to other sites.

Yuri worked on the Russian laptop #1, creating a log file of its data on a floppy disk for subsequent downlink via the REGUL-Packet system and analysis at MCC-M.

Malenchenko also continued his current (seventh) session of the Russian Uragan (“hurricane”) earth imaging program (GFI-8), using the Kodak DCS760 digital still camera with 800-mm lens where cloud cover did not exceed 30%.  [Targets included the Jarkuduk region and RGO glacier, the Nurek hydro-electric power station, the Toktogul water reservoir, the northern coast of the Cheleken Peninsula, open pits at Temirtau and Karaganda, and the flood in Tuva (Khovu-Aksy region).  The images, in TIF format, were again stored on the removable hard disk drive (HDD) #4 of the French EGE-2 laptop.]

For the Diatomeya ocean observations program, Yuri completed another photography session, using the DVCAM 150 digital video camera and the Nikon-F5 electronic still camera with 80-mm lens from SM windows #8 and #7.  [Today’s observations presented an opportunity to investigate bioproductive waters in the Sea, Black Sea and the Sea of Azov, a repeat of observations done on 8/5.]

Ed Lu performed the regular daily routine tasks of SOZh life support system maintenance, inventory management system (IMS) delta file preparation, and status checkup of the autonomous Lab payloads.  Yuri Malenchenko tagged up with the ground on the weekly IMS status update.  [The telecon today focused on a discarded USIL (= IELK, individual equipment and liner kit) soft cover, the missing barcode number of a KRP food container, an update on BRPK condensate separation units aboard ISS, and a review of the IMS loading plan for Progress-259/11P.]

Both crewmembers worked out in their daily 2.5-h program of physical exercise, on TVIS treadmill and RED expander. 

Ed did another transfer of data files from the exercise equipment to the MEC (medical equipment computer) and worked on the RED.  [Objectives of his IFM (in-flight maintenance) were to (1) relieve compression in the canister and (2) install shims to maintain decompression to achieve nominal operations including full cord retraction.  The lowering of compression was aimed at solving the cord retraction issue of the current canister (i.e., pulley cords not fully retracting), and increase performance on the cans.]

At 12:25pm EDT, Lu set up and conducted a ham radio session with students at Soar Valley College in Leicester, England.  [Soar Valley is a secondary school with 1200 students.  It has established an amateur radio, satellite and computing club specifically to promote interest in technology, space studies and amateur radio training with novice instruction.  The ISS crew is always glad to help with this.]

The cooling anomaly encountered yesterday by Ed Lu with EMU/spacesuit #3005 continues to be analyzed on the ground.  [EMU 3005 had been brought back to life after almost a year of on-orbit “hibernation”.  The crew’s schedule currently provides for a 15 min time block on 8/12 for stowing any ancillary EMU items that may still be unstowed.  Next on the EVA agenda is return to troubleshooting Ed’s original EMU, #3013.  This is currently scheduled for 8/15.  It will hopefully isolate the cause of its cooling problem and exonerate the LCVG (liquid cooling ventilation garment) and SCU-1 (servicing & cooling umbilical #1) water condition.  And no  Ed will not have to wiggle back into the suit.]

Because of the EMU activities, yesterday’s planned EPO (Educational Payload Operations) had to be deferred.  The educational demonstration of EVA and IVA tools is now being rescheduled for next week, probably in place of the originally planned EPO Newton’s Law demo.

MCC-H is continuing troubleshooting in support of the SSC4 (station support computer #4) laptop.  [From info called down by Ed Lu regarding possible application and/or network problems with SSC4, additional troubleshooting steps are being recommended by ground specialists, to determine if the difficulties are a minor software config anomaly or lingering hardware problems that have been seen in the recent past.]

Starting 8/5 and running through 8/10, MCC-H and POC (Payload Operations Center) are continuing a Ku-band 150 MBPS (megabytes per second) evaluation test, during which the IP Phone and OCA transfer capabilities are temporarily out of service.  [Currently, MCC-H and POC use the 50 Mbps capability of Ku-band.  The purpose of this test is to demonstrate that the ISS Ku-band 150 Mbps and 75 Mbps can be effectively utilized for future day-to-day operations.]

MCC-M/TsUP has no command capability from 8/7 through 8/9 (Saturday) due to maintenance at the Shelkovo ground station and unavailability of the Molniya satellite.  Voice capability is not affected during this period.  [RGS Shelkovo telemetry equipment will be operational a day earlier, on 8/9, but command equipment will not be ready until 8/10 (Sunday).]

Also on 8/9, there will be a 3-hr. outage of Ku-band downlink capability at MCC-H, to support installation of new power (high-rate data switch unavailable), starting at 11:30am EDT.

At regular intervals, the station’s cabin atmosphere is refreshed with gaseous oxygen (O2) from Progress 10P O2 tanks as long as its supply lasts (originally 50 kg/110 lbs), whenever ppO2 (O2 partial pressure) decreases to 160 mmHg.  Afterwards, O2 repress will also be performed from Progress 11P.  The refresh was not required today.  [During the Progress-supplied repress period, the Elektron oxygen generator is off, to conserve its service life, and the repress valve cycling, by the crew, relies on the U.S. MCA (major constituents analyzer) or CSA-CP (compound specific analyzer-combustion products) instruments to measure the 10mmHg pressure difference.]

The VOA (Volatile Organics Analyzer) is currently failed.

Today’s CEO (crew earth observation) targets, no longer limited in the current LVLH attitude and including the targets of the Lewis & Clark 200-year memorial locations, were Betsiboka River delta, Madagascar (Dynamic event.  Sunglint opportunity, a touch right of track.  This largest estuary on Madagascar’s north coast has experienced very fast sedimentation due to deforestation inland.  A new island in the estuary was documented last year.  Images for this year are needed to document the progress of mangrove colonization across the new island. Growth of the previous “new” has been well documented in handheld images.  The CEO image would permit publication of this interesting story), Congo-Zimbabwe Biomass Burning (fires south of the Congo making the news.  Looking left and right of track for fires and any regional smoke pall), Egypt – Valley of the Kings (looking a touch right of track, immediately west of the Nile where the Nile Valley floor narrows), Statue of Zeus at Olympia (nadir pass.  Looking on the west coast of the Peloponnesos peninsula in the largest valley entering the Med), Athens, Greece (pointing a touch right), Istanbul, Turkey (pointing a touch right), Brasilia, Brazil (nadir pass), Berlin, Germany (pointing a touch left for this large, tree-dominated city.  The heat wave in Europe associated with the clear weather is breaking records.  Air flow from the Sahara is partly responsible), London, England, Great Britain (nadir pass), Chicago, Illinois (nadir pass), Fort Mandan, ND (LEWIS & CLARK SITE:  The site is near the present town of Stanton. Look a touch left on the Missouri River, on the south side of the large reservoir, Lake Sacagawea.  Stanton is on the west bank of the Missouri near where the river turns sharply from south to east.  The expedition spent the winter of 1804-5 here, and Sacagawea and her husband joined the group), and Fort Manuel, SD (pointing a touch right on the river).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:03pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off (due to low humidity); SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.3; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.5.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 734.98; temperature (deg C) — 22.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — 4.6.
  • U.S. Lab:  Pressure (mmHg) — 737.24; temperature (deg C) — 23.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 737.34; temperature (deg C) — 22.7; shell heater temp (deg C) — 21.6, ppO2 (mmHg) — 161.7; ppCO2 (mmHg) — 3.6.
  • PMA-1:  Shell heater temp (deg C) — 23.2
  • PMA-2:  Shell heater temp (deg C) — 19.8.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3646 kg (8038 lb) as of 7/31  [SM(774) + FGB(2872) + Progress M(0) +Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (blind, dual-angle mode with drag reduction, non solar-tracking).
  • SM batteries:  Battery #5 is off line(Capacity Restoration mode, ROM); all other batteries (7) are in “Partial Charge” mode.  Batteries #1and #3 showing degraded capacity. 
  • FGB batteries:  Battery #1 is off line (Capacity Restoration mode, ROM); battery #4 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.
  • Solar Beta angle:  1.3 deg (magnitude increasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 00:31am EDT [= epoch]):

  • Mean altitude — 384.7 km
  • Apogee — 390.2 km
  • Perigee — 379.2 km
  • Period — 92.25 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.000817
  • Orbits per 24-hr. day — 15.61
  • Solar Beta Angle — 1.3 deg (magnitude increasing)
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26912
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.